# Distributed Certifiably Correct Range-Aided SLAM

### v1

@@ IV-B3 Discussion
Results demonstrate comparable ATE performance between DCORA and CORA. This is expected, as these algorithms are (mainly) differentiated by the order of their optimization procedures (i.e., first- vs. second-order). The order of optimization does not affect the optimality of the SDP; thus, the certified solution (mathematically) should remain identical. However, RBCD has a sublinear convergence rate, meaning its estimates are not as precise as those provided by CORA’s second-order trust-region method. As a result, DCORA requires a looser certification tolerance to verify optimality, which accounts for the small differences in $f^{\ast}_{\text{SDP}}$ and ATE on MR.CLAM.
@@ IV-C3 Discussion
As observed in Fig. 3, convergence rate and solution precision are predominantly affected by the number of agents. This trend is most clear on the modified smallGrid3D dataset, where fewer RBCD iterations are required to converge to a lower Riemannian gradient norm, thus suggesting faster convergence to more precisely estimated critical points. On the modified sphere2500 dataset, an exception to this trend is observed when the range measurement probability is 1.0. Specifically, the eight-agent system demonstrates improved convergence and solution precision compared to the four-agent system. We conjecture that the increase in range measurement density compensates for the loss in connectivity in each agent’s underlying pose graph, as intra-agent relative pose loop closures are redistributed as inter-agent loop closures. Observations on SDP tightness further support this conjecture, as increasing range measurement density improves tightness as the number of agents increases.
@@ V Conclusion
In this work, we presented DCORA, the first distributed certifiably correct algorithm for range-aided simultaneous localization and mapping (RA-SLAM). DCORA generalizes the Riemannian Staircase method established for distributed certifiably correct pose graph optimization to the RA-SLAM problem and demonstrates comparable solution quality to its centralized counterpart, CORA, on real-world datasets. The structure of the RA-SLAM problem in the distributed setting was studied using parametrically varied synthetic datasets, revealing the number of agents to be the predominant factor affecting DCORA’s convergence rate and solution precision. Results also showed that graph connectivity influences SDP tightness, where additional range measurements can aid solution quality for systems distributing relative pose measurements across a large number of agents. In future work, we aim to expand the real-world evaluation of DCORA across larger multi-agent systems under various noise regimes. We also plan on studying DCORA’s runtime performance with a fully distributed implementation. Of broader importance to DCORA’s method class, efforts to significantly enhance the speed, scalability, and communication efficiency of Riemannian block coordinate descent by incorporating (at least partial) second-order information into its design would be a substantial contribution towards enabling efficient estimation and control methods for multi-agent systems in general. This, in combination with identifying sufficient conditions to ensure that DCORA’s (and CORA’s) SDP relaxation remains exact under bounded noise for the general RA-SLAM problem, are directions for future work.

### v2

@@ IV-B3 Discussion
Results demonstrate comparable ATE performance between DCORA and CORA. This is expected, as these algorithms are (mainly) differentiated by the order of their optimization procedures (i.e., first- vs. second-order). The order of optimization does not affect the optimality of the SDP; thus, the certified solution (mathematically) should remain identical. However, RBCD has a sublinear convergence rate, meaning its estimates are not as precise as those provided by CORA’s second-order trust-region method. As a result, DCORA requires a looser certification tolerance to verify optimality, which accounts for the small differences in $f^{\ast}_{\text{SDP}}$ and ATE.
@@ IV-C3 Discussion
As observed in Fig. 3, convergence rate and solution precision are predominantly affected by the number of agents. This trend is most clear on the modified smallGrid3D dataset, where fewer RBCD iterations are required to converge to a lower Riemannian gradient norm, thus suggesting faster convergence to more precisely estimated critical points. For the modified sphere2500 dataset, an exception to this trend is observed when the range measurement probability is 1.0. Specifically, the eight-agent system demonstrates improved convergence and solution precision compared to the four-agent system. We conjecture that the increase in range measurement density compensates for the loss in connectivity in each agent’s underlying pose graph, as intra-agent relative pose loop closures are redistributed as inter-agent loop closures. Observations on SDP tightness further support this conjecture, as increasing range measurement density improves tightness as the number of agents increases.
@@ V Conclusion
In this work, we presented DCORA, the first distributed certifiably correct algorithm for range-aided simultaneous localization and mapping (RA-SLAM). DCORA generalizes the Riemannian Staircase method established for distributed certifiably correct pose graph optimization to the RA-SLAM problem and demonstrates comparable solution quality to its centralized counterpart, CORA, on real-world datasets. The structure of the RA-SLAM problem in the distributed setting was studied using parametrically varied synthetic datasets, revealing the number of agents to be the predominant factor affecting DCORA’s convergence rate and solution precision. Results also showed that graph connectivity influences SDP tightness, where additional range measurements can aid solution quality for systems distributing relative pose measurements across a large number of agents. In future work, we aim to expand the real-world evaluation of DCORA across larger multi-agent systems under various noise regimes. We also plan on studying DCORA’s runtime performance with a fully distributed implementation. Of broader importance to DCORA’s method class, efforts to significantly enhance the speed, scalability, and communication efficiency of Riemannian block coordinate descent by incorporating (at least partial) second-order information into its design would be a substantial contribution towards enabling efficient estimation and control methods for multi-agent systems in general. This, in addition to identifying sufficient conditions to ensure that DCORA’s (and CORA’s) SDP relaxation remains exact under bounded noise for the general RA-SLAM problem, are directions for future work.